# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

### v1

@@ V-C Discussion on Post-computation Communication
In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional client-to-client communications following the execution of the PIC protocol. For instance, consider a ridesharing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$. After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up. On the other hand, the drivers also need to share their identities and locations with the matched passengers. Inevitably, a party has to sacrifice their privacy to the matched parties.
The post-computation communication may also have privacy implications for other clients not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the clients in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
@@ VI-A Inadequacy of Existing Randomizers
In the PIC model, each user first sanitizes their data using an $\epsilon$-LDP randomizer. The design of the randomizer significantly impacts the utility of the tasks. While the PIC model can be seen as an extension of the shuffle model, it has unique characteristics that render the existing LDP randomizers commonly used in the shuffle model inadequate.
The main discrepancy between the shuffle model and the PIC model is that the former emphasizes statistical utility, whereas the latter focuses on the utility of each report. The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic. In the literature, utility is often measured by the expected square error (i.e., the variance) of the estimation:
$\mathbb{E}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\textsf{Var}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\tilde{f}(T)]$
@@
To see the problem, we first examine a class of LDP randomizers [64, 84, 58, 50]) that work by sampling a few dimensions from $[d]$. Each user submits an incomplete report that contains only the sampled dimensions (with added noise) in their local data. This is an effective strategy for statistical estimation, but will totally ruin the utility of a single report in PIC as unsampled dimensions are missing. Clearly, they are unsuitable for use in the PIC model.
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [27], PlanarLaplace [4], PrivUnit [18], and PrivUnitG mechanisms [8]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [9], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -B).
@@ VI-B Randomizer Design
We now introduce an asymptotic optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -B), but the mechanism can be generalized to other Minkowski distances.
Without loss of generality, we assume the user data domain to be a $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$. Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g. gradient vector, set-valued data, and location data). We also denote an $\ell_{2}$-bounded hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$ as follows:
@@
Figure 4: Illustrative probability design of Minkowski response mechanism over two dimensions with cap area radius $r$, given different inputs $x_{i}$ in (a) and (b) respectively.
@@
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 5 in Appendix -B), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
@@ VI-C Analysis of Minkowski Response
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{X}$, we have $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -B.
@@
Given Theorems VI.2 and VI.3, we observe that as $n\rightarrow\infty$, the error upper bound $\tilde{O}(\frac{1}{n^{2/(d+2)}})$ of the Minkowski Response randomizer is the same as the error lower bound of the PIC model (the $\tilde{O}$ notation suppresses factors about $\epsilon_{c},\delta$ for simplicity). This means that the utility of the Minkowski Response is asymptotically optimal. In contrast, existing LDP randomizers with privacy amplification in the PIC model will suffer the single report error of $\tilde{O}(\frac{d}{n^{2/(3d)}})$ or $\tilde{O}(\frac{d}{\log^{2}n})$, whose denominator grows much slower than $n^{2/(d+2)}$ when $d>1$ and $n$ is relatively large. In practice, $n$ does not need to be very large; in our experiments, Minkowski Response outperforms existing LDP randomizers in PIC model once $n$ reaches the order of $10^{2}$. When comparing the PIC model to local DP model, the gap is undoubtedly larger, as numerical methods of privacy amplification [56, 85] will always give a local budget $\epsilon$ that is larger than the global privacy $\epsilon_{c}$ as long as $n\geq 2$, creating an extra gap in between the local and PIC models of these LDP randomizers. Actually, any randomizers in the LDP model must endures an error of $\Omega({d}/{\epsilon_{c}^{2}})$ for single report when $\epsilon_{c}\leq O(1)$ [18, 25].
@@ VII Experimental Evaluation
We evaluate the utility and efficiency performances of proposed PIC model in three representative individual computation tasks: spatial crowdsourcing, location-based nearest neighbors, and federated learning with incentives, aim to answer the following research questions:
  • RQ1.
    How much do our proposals outperform existing approaches in accuracy/utility?
  • RQ2.
    What the computational/communication costs of our model for supporting security functionalities?
TABLE II: Mean $\ell_{2}$-error (and running time) comparison of local $\epsilon$-LDP randomizers on location domain $[-1,1]\times[-1,1]$. All results are the expected value of $1000$ repeated experiments.
randomizer $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$ $\epsilon=3.0$ $\epsilon=5.0$ $\epsilon=8.0$ $\epsilon=10.0$
PrivUnit [16] 8.94 (49us) 4.68 (42us) 2.25 (61us) 1.44 (63us) 0.81 (113us) 0.32 (305us) 0.18 (868us)
PrivUnitG[8] 8.73 (76ms) 4.63 (75ms) 2.27 (77ms) 1.51 (76ms) 0.96 (74ms) 0.63 (75ms) 0.53 (81ms)
Laplace [28] 12.97 (0.1us) 6.56 (0.1us) 3.27 (0.1us) 2.13 (0.1us) 1.30 (0.1us) 0.81 (0.1us) 0.64 (0.1us)
PlanarLaplace [4] 11.17 (12us) 5.63 (11us) 2.84 (11us) 1.88 (11us) 1.14 (11us) 0.71 (12us) 0.56 (12us)
Staircase[35] 13.19 (49us) 6.40 (48us) 3.13 (46us) 2.01 (48us) 1.05 (44us) 0.46 (51us) 0.28 (50us)
SquareWave[57] 11.87 (1.9us) 5.72 (2.5us) 2.65 (1.9us) 1.68 (2.2us) 0.92 (1.7us) 0.53 (2.2us) 0.42 (2.1us)
MinkowskiResponse 10.42 (0.7us) 4.50 (0.6us) 1.78 (0.8us) 0.98 (0.8us) 0.39 (0.6us) 0.14 (0.7us) 0.074 (0.8us)
@@ VII-A Spatial Crowdsourcing
Datasets and settings. In the experiments for spatial crowdsourcing, we use use two real datasets, GMission dataset [20] and EverySender dataset [79], for scientific simulation and campus-based micro-task completion respectively. The number of users/workers, location domain range, and serving radius of workers about these two datasets are summarized in Table III (more information are presented in Appendix -F). In our experiments, for compliance with data randomizers (see following parts), we rescale the location domain range to $[-1,1]\times[-1,1]$ and rescale the corresponding serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4.
TABLE III: Summary statistics of spatial crowdsourcing datasets.
@@
Cryptography implementation. Follow the protocol in Section V, we implement the PIC model using Elliptical Curve Diffie-Hellman key agreement (ECDHE [11]) for establishing shared secret between client/server and matched user/worker, meanwhile using AES encryption [68] and HMAC signature [14] for secure point-to-point communication after key agreement.
Data randomizers and server-side algorithms. We use the Minkowski response for location randomization. As comparison, in the local model of DP (e.g., in [83, 77, 81]), we compare with existing mechanisms including Laplace [28], Staircase [35], PlanarLaplace with geo-indistinguishability [4], SquareWave [57], PrivUnit [16] and its Gaussian variant PrivUnitG in [8]. On the server side, we use either minimum weighted full matching [51] or maximum matching [45] over bipartite graph as the user/worker matching algorithm based on (noisy) locations, to obtain a match $M:G_{a}\times G_{b}\mapsto\mathbb{R}$. Specifically, the minimum weighted full matching aims to minimize the overall travelling costs between users and works over the (noisy) fully-connected bipartite graph; the maximum matching aims to maximize the number of successfully matched user/worker pairs over the (noisy) partially-connected bipartite graph, where user/worker pairs with distance surpassing $0.4$ (i.e. the serving radius) are deemed as not connected.
Performance metrics. We evaluate the performance of spatial crowdsourcing in terms of travel costs, successful matching ratio, and running time. The travel costs are the total actual Euclidean distance between all pairing matched user/workers:
$\text{travel costs}=\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
@@
$\text{success ratio}=\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
We evaluate both the client-side and server-side running time of our protocol on a laptop computer embedded with Intel i5555-8250U CPU @1.6GHz and $8$GB memory.
Figure 5: Expected $\ell_{2}$ error of reported locations.
Performance of location reports. Before exploring spatial crowdsourcing tasks, we first evaluate our proposed Minkowski response mechanism against existing mechanisms, as presented in Table II. This evaluation focuses on errors and running times associated with reporting locations under local $\epsilon$-DP constraints. We quantify the error using the expected $\ell_{2}$ distance between each true and noisy location pair. Notably, the Minkowski response showcases an $\ell_{2}$ error comparable to state-of-the-art (SOTA) mechanisms when $\epsilon\leq 1.0$. Its superiority is evident when $\epsilon\geq 2.0$, displaying a significantly reduced error. This supports the theoretical assertions made in Section VI, underscoring the importance of the Minkowski response for the shuffle model, over other existing randomizers. Additionally, the Minkowski response is highly efficient, demanding less than $1$us on the user end. This speed is nearly equivalent to merely adding Laplace noises and is roughly $50$x faster than the previously SOTA PrivUnit.
We further contrast the utility performances of the mechanisms in the local model with those in the shuffle model. Privacy amplification in the shuffle model is achieved using the latest analyses from [32]. As illustrated in Figure 5, the shuffle model presents a significantly reduced location reporting error compared to the local model. This disparity becomes more pronounced as the number of users increases (notably, the EverySender dataset comprises more users than GMission). The introduction of an apt randomizer—such as the Minkowski response—can further elevate utility. This is evident when comparing the Minkowski response and PrivUnit, both of which operate within the shuffle model.
Performance of minimum weight matching. Figure 6 presents the expected travel costs associated with minimum weighted full matching. For a comprehensive overview of the experimental results involving more competitors, readers are directed to -F. Notably, the Squarewave mechanism outperforms PrivUnit in matchings, prompting its primary representation in the shuffle model discussions. The equivalent shuffle model yields a cost reduction of approximately $50\%$, with the savings escalating as the number of users increases. The Minkowski response consistently exhibits the lowest costs across all scenarios. Together, they result in a substantial reduction of about $75\%$ when compared to existing approaches.
Figure 6: Travel costs of minimum weighted matching.
Performance of maximum matching. Figure 7 illustrates the success ratios for maximum matching. The equivalent shuffle model shows a relative improvement of approximately $200\%$ in the success ratio. The Minkowski response consistently yields higher ratios across all scenarios. When the user count becomes relatively substantial (for instance, a few thousands as in EverySender), our results indicate that spatial crowdsourcing maximum matching, with a reasonable level of privacy protection (specifically, $\epsilon\leq 3$), can be made practical, achieving a success ratio exceeding $90\%$.
Figure 7: Success ratios of maximum matching.
TABLE IV: Running time and communication overheads of spatial crowdsourcing in the PIC model.
User-side Procedures Time Communication Server-side Procedures Time Communication
location randomization 0.7us - decrypt one message 2.1ms -
encrypt information 2.9ms - min-weight match (GMission) 31ms 1.2MB
send message to shuffler - 1.3KB maximum match (GMission) 15ms 1.2MB
retrieve matching result - 1.8KB min-weight match (EverySender) 79ms 4.3MB
key agreement with worker 2.6ms - maximum match (EverySender) 63ms 4.3MB
Running time and communication overheads. We present the computation and communication overheads of the PIC model in Table IV. When considering multiple procedures on the user side, the total running time for each user is only a few milliseconds (with the data randomization time being negligible), and the total communication overhead is under $4$KB. On the server side, the running time of the matching algorithm on noisy plaintext is consistent with non-private settings, scaling with the number of users/workers and taking dozens of milliseconds. The additional decryption/encryption runtime is capped at several seconds, mirroring the costs of HTTPS decryption/encryption. Specifically, it utilizes ECDHE for key agreement and AES encryption as with HTTPS on TLS 1.3 [67] for client-server communication. In essence, the running times on both the client and server sides of the PIC model closely resemble those in non-private settings that use HTTPS communication (for a fair comparison).

### v2

@@ V-C Discussion on Post-computation Communication
In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional user-to-user communications following the execution of the PIC protocol. For instance, consider a taxi-hailing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$. After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up. On the other hand, the drivers also need to share their identities and locations with the matched passengers. Inevitably, a party has to sacrifice their privacy to the matched parties.
The post-computation communication may also have privacy implications for other users not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the users in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
@@ VI-A Inadequacy of Existing Randomizers
In PIC, each user first sanitizes their data using an $\epsilon$-LDP randomizer. The design of the randomizer significantly impacts the utility of the tasks. While the PIC model can be seen as an extension of the shuffle model, it has unique characteristics that render the existing LDP randomizers commonly used in the shuffle model inadequate.
The main discrepancy between the shuffle model and the PIC model is that the former emphasizes statistical utility, whereas the latter focuses on the utility of each report. The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic. In the literature, utility is often measured by the expected square error (i.e., the variance) bound of the estimation:
$\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\mathbb{E}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\textsf{Var}[\tilde{f}(T)]$
@@
To understand the problem, we first examine a class of LDP randomizers [65, 86, 59, 51] that operate by sampling a few dimensions from $[d]$. Each user submits an incomplete report that contains only the sampled dimensions (with added noise) from their local data. On the positive side, this strategy reduces the amount of noise added to the sampled dimensions. On the negative side, the unsampled dimensions are missing. In statistical estimation tasks, the estimation is made using all reports, each covering some dimensions. Therefore, the incompleteness of a single report is less important, and better utility can be achieved. However, in the PIC model, where the focus shifts to the error of individual reports, this strategy may lead to worse results.
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [30], PlanarLaplace [2], PrivUnit [12], and PrivUnitG mechanisms [6]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [7], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -D).
@@ VI-B Randomizer Design
We now introduce an asymptotically optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -D), but the mechanism can be generalized to other Minkowski distances.
Without loss of generality, we assume the user data domain to be an $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$. Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g. gradient vector, set-valued data, and location data). We also denote an $\ell_{2}$-bounded hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$ as follows:
@@
Figure 4: The probability design of Minkowski response mechanism with a radius $r$. Illustrated are three inputs $x_{1},x_{2}$ and $x_{3}$, along with their respective cap areas.
@@
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 6 in Appendix -D), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
@@ VI-C Analysis of Minkowski Response
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$, we have:
$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -D.
@@
We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters. If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem VI.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem VI.2. This implies that the utility of the Minkowski Response is asymptotically optimal. In contrast, using existing LDP randomizers in the PIC model results in a larger single report error of $\tilde{O}\left(\frac{d}{n^{2/(3d)}}\right)$ or $\tilde{O}\left(\frac{d}{\log^{2}n}\right)$ (note that $d>1$ and $n$ is often not small). Although asymptotic notations describe behavior as $n\rightarrow\infty$, in practice, the utility advantage of the Minkowski Response becomes noticeable without $n$ being very large: in our experiments, the Minkowski Response randomizer outperforms existing LDP randomizers in the PIC model once $n$ reaches the order of $10^{2}$. If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
@@ VII Experimental Evaluation
We evaluate the efficacy of our PIC protocol and Minkowski randomizer. We compare the utility of our proposal against state-of-the-art works in the context of three representative individual computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives.
@@ VII-A Spatial Crowdsourcing
Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [82] for campus-based micro-task completion. Details about the two datasets are summarized in Table II, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix -F). We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
TABLE II: Summary statistics of spatial crowdsourcing datasets.
@@
LDP randomizers We use the Minkowski response for location randomization. As a comparison, in the local model of DP (e.g., in [85, 80, 84]), we compare with existing mechanisms including Laplace [31], Staircase [38], PlanarLaplace with geo-indistinguishability [2], SquareWave [58], PrivUnit [12] and its Gaussian variant PrivUnitG in [6].
Server-side algorithms Two commonly used server-side algorithms are evaluated as concrete tasks: minimum weighted full matching [52] and maximum matching [47]. The two algorithms have different optimization objectives, thus later we will show the results for task-specific utility for each of them in addition to single report errors. The minimum weighted full matching aims to minimize the overall traveling costs between users and workers. The maximum matching aims to maximize the number of successfully matched user/worker pairs, where users outside the workers’ serving radius ($0.4$) are deemed unreachable.
Experimental results We first present the single report errors, quantified by the expected $\ell_{2}$ distance between the reported location and the true location. In Fig. 5 (a) and (b), we compare the single report errors within the LDP model. The Minkowski randomizer’s error is on par with other state-of-the-art mechanisms when $\epsilon\leq 1.0$ and significantly lower when $\epsilon\geq 2.0$. Fig. 5 (c) and (d) illustrate the single report errors in the PIC model. For both datasets, the Minkowski randomizer performs better, with its utility advantage increasing as the global privacy budget $\epsilon_{c}$ grows. Additionally, it is evident that the error is generally higher for all randomizers in the LDP model compared to the PIC model. This discrepancy is due to the lack of privacy amplification in the LDP model, highlighting the benefits of employing the PIC model.
Figure 5: Expected $\ell_{2}$ distances of reported locations to true locations on GMission and EverySender dataset.
Next, we compare the utility when using the minimum weight matching algorithm, as shown in Fig. 6. Task-specific utility is evaluated by the total travel costs, which is the sum of the actual Euclidean distances between all matched users/workers:
$\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
@@
Figure 6: Travel costs of minimum weighted matching on GMission and EverySender dataset.
Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, task-specific utility is assessed by the successful matching ratio:
$\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
For both datasets, the matching ratio over clear data is 100%. The figure demonstrates that the PIC model enhances the matching ratio for all randomizers due to privacy amplification effects. The Minkowski randomizer in the PIC model significantly outperforms the others and the matching ratio approaches an acceptable level for practical use with a reasonable degree of privacy protection.
Figure 7: Success ratios of maximum matching in spatial crowdsourcing on GMission and EverySender dataset.